A Note About Invariant Polynomial Transformations of Integer Sequences

Leonid Bedratyuk Department of Applied Mathematics Khmelnitskiy National University Instituts’ka 11 Khmelnitskiy, 29016 Ukraine

                      leonid.uk@gmail.com

Abstract

We present an algorithm to find invariant poynomial transformations of integer sequences, using an approach based on classical invariant theory.

1 Introduction

Let A = (a_{n})_{n≥0} be an integer sequence. A sequence F(A) = (b_{n} = f_{n}(a_{0}, a_{1}, …, a_{m}))_{n≥0}, where f_{n} ∈ ℤ[x_{0}, x_{1}, …, x_{m}] and m ≥ n, is called a polynomial transformation of the sequence A. In the sequel, only the polynomial transformations are considered. The composition F ∘ G := F(G(A)) of the two transformations F and G can be defined in a natural way. A transformation G is called the inverse transformation of F, and it is denoted by F^{−1}, if for every sequence A we have F(G(A)) = A. A transformation F is called G-invariant if for every sequence A we have F(G(A)) = F(A).

For instance, it is well known (see Layman [1] or Spivey and Steil [2]) that the Hankel transformation H is B_{μ}-invariant. Here, for μ ∈ ℚ,

B_{μ}(A) = (b_{n} = ∑_{i=0}^{n} \binom{n}{i}a_{i}μ^{n−i})_{n≥0} ,

denotes the μ-binomial transformation and H(A) = (h_{n})_{n≥0}, where h_{n} is the determinant of the Hankel matrix for the elements a_{0}, a_{1}, …, a_{2n}:

h_{n} = |\begin{matrix}a_{0} & a_{1} & a_{2} & ⋯ & a_{n} \\ a_{1} & a_{2} & a_{3} & ⋯ & a_{n+1} \\ ⋮ & ⋮ & ⋮ & ... & ⋮ \\ a_{n−1} & a_{n} & a_{n−1} & ⋯ & a_{2n−1} \\ a_{n} & a_{n+1} & a_{n+2} & ⋯ & a_{2n} \end{matrix}| .

This determinant arises in classical invariant theory as the catalecticant of a binary form, see Grace and Young [4, p.232]. It was first introduced by Sylvester [5]. Also, the transformation B_{μ} appears in Hilbert’s book [7, p. 25]. We can prove that the Hankel transformation is B_{μ}- invariant via classical invariant theory, as follows. Write ∂_{i} for the partial derivative ∂/∂a_{i} and let D be the differential operator:

D = a_{0}∂_{1} + 2a_{1}∂_{2} + 3a_{2}∂_{3} + ⋯ + 2na_{2n−1}∂_{2n}.

Define

h′_{n} = |\begin{matrix}b_{0} & b_{1} & b_{2} & ⋯ & b_{n} \\ b_{1} & b_{2} & b_{3} & ⋯ & b_{n+1} \\ ⋮ & ⋮ & ⋮ & ... & ⋮ \\ b_{n−1} & b_{n} & b_{n−1} & ⋯ & b_{2n−1} \\ b_{n} & b_{n+1} & b_{n+2} & ⋯ & b_{2n} \end{matrix}| .

Then, as in Lie [6], we have

h′_{n} = h_{n} + D(h_{n})μ + D^{2}(h_{n})\frac{μ^{2}}{2!} + ⋯ + D^{i}(h_{n})\frac{μ^{i}}{i!} + · · · .

By applying the determinant derivative rule we get, after some calculation, that D(h_{n}) = 0 for all n. Therefore h′_{n} = h_{n}. This condition is exactly equivalent to the B_{μ}-invariance of the Hankel transformation.

This example motivates us to consider the following two general problems:

Problem 1. For a fixed transformation F, find all F-invariant transformations.

Problem 2. For a fixed transformation F, find a transformation G such that F is a Ginvariant transformation.

The aim of this paper is to develop an effective method to solve these two problems for some special kinds of transformations. The method is inspired by results in classical invariant theory and the theory of locally nilpotent derivations.

In the next section we introduce exponential transformations and then prove that for such transformations Problem 1 is always solvable.

In Section 2 we also give a short introduction to the theory of locally nilpotent derivations and offer algorithms to solve Problems 1 and 2 for special classes of transformations.

In Section 3 we give another proof that the Hankel transformation is B_{μ}-invariant and introduce several new B_{μ}-invariant transformations. All of these transformations come from classical invariant theory. Further, we describe all B_{μ}-invariant polynomial transformations in terms of derivations.

In Section 4 we give some examples to illustrate the theory.

2 Derivations and automorphisms

Let R = ℤ[x_{0}, x_{1}, …] be the polynomial ring in countably many variables and let φ : R → R be a ring homomorphism. Such a map φ is uniquely determined by the sequence of polynomials (φ(x_{n}) | n = 0, 1, …). To any polynomial map φ and an integer sequence (a_{n})_{n≥0} we assign the transformation (φ(a_{n}))_{n≥0}. A polynomial map φ is said to be a polynomial automorphism if there is a polynomial map ψ such that φ(ψ(x_{n})) = x_{n} for all n.

Denote by ℤ[x_{0}, x_{1}, …, x_{m}]^{φ} the algebra of φ-invariants:

ℤ[x_{0}, x_{1}, …, x_{m}]^{φ} := {f ∈ ℤ[x_{0}, x_{1}, …, x_{m}] | f(φ(x_{0}), φ(x_{1}), …, φ(x_{m})) = f(x_{0}, x_{1}, …, x_{m})} .

The following theorem will be our main computing tool in finding invariant polynomial transformations.

Theorem 1. Let φ be a polynomial map and let F(A) = (φ(a_{n}))_{n≥0} be the corresponding integer transformation. Then the transformation

G(A) = (g_{n}(a_{0}, a_{1}, …, a_{m}))_{n≥0} ,

is F-invariant if and only if g_{n}(x_{0}, a_{1}, …, x_{m}) ∈ ℤ[x_{0}, x_{1}, …, x_{m}]^{φ}.

The proof follows immediately from the definitions above.

In general, the problem of finding the algebras of φ-invariants is difficult. But we will show that when φ is an exponential automorphism this problem can be reduced to the calculation of the kernel of a derivation.

A derivation of the algebra ℤ[x_{0}, x_{1}, …, x_{n}] is a linear map D satisfying the Leibniz rule:

D(f_{1} f_{2}) = D(f_{1})f_{2} + f_{1}D(f_{2}), for all f_{1}, f_{2} ∈ ℤ[x_{0}, x_{1}, …, x_{n}].

A derivation D is called locally nilpotent if for every f ∈ ℤ[x_{0}, x_{1}, …, x_{n}] there is an r ∈ ℕ such that D^{r}(f) = 0. The subalgebra

ker D := {f ∈ ℤ[x_{0}, x_{1}, …, x_{n}]|D(f) = 0} ,

is called the kernel of the derivation D.

Any derivation D is completely determined by the elements D(x_{i}). A derivation D is called linear if D(x_{i}) is a linear form. A linear locally nilpotent derivation is called a Weitzenböck derivation. The Weitzenböck derivation defined by D(x_{0}) = 0, D(x_{i}) = ix_{i−1} is called the basic Weitzenböck derivation. There exists an isomorphism between the kernel ker D and the algebra of covariants of a binary form, a major object of research in classical invariant theory during the 19th century. Here are a few examples of covariants: the discriminant, the resultant, the Jacobian, the Hessian, the catalectiant and the transvectant. The following theorem gives a description of the algebra ker D.

Theorem 2. The kernel of the basic Weitzenböck derivation D of ℚ[x_{0}, x_{1}, …, x_{n}] is a finitely generated algebra and

ker D = ℚ[z_{2}, z_{3}, …, z_{n}][x_{0}, x_{0}^{−1}] ∩ ℚ[x_{0}, x_{1}, …, x_{n}],

where

z_{k} = ∑_{i=0}^{k−2}(−1)^{i}\binom{k}{i}x_{k−i}x_{1}^{i}x_{0}^{k−i−1} + (k − 1)(−1)^{k+1}x_{1}^{k}.

Theorem 2 is a classical result due to Cayley, see Glenn [3, p. 164]. One can also find more modern proofs in Nowicki [8] and van den Essen [9].

How are we to find the kernel of an arbitrary linear locally nilpotent derivation D? Let us consider the vector space (over ℚ) X_{n} = ⟨x_{0}, x_{1}, …, x_{n}⟩. Suppose that there exists an isomorphism Ψ : X_{n} → X_{n} such that ΨD = DΨ. This implies that ker D = Ψ (ker D) , i.e.,

ker D = ℚ[Ψ(z_{2}), Ψ(z_{3}), …, Ψ(z_{n})][Ψ(x_{0}), Ψ(x_{0})^{−1}] ∩ ℚ[x_{0}, x_{1}, …, x_{n}].

Such an isomorphism Ψ is called a (D, D)-intertwining isomorphism. Therefore, to describe the kernel of an arbitrary Weitzenbök derivation D it is enough to know the explicit form of any (D, D)-intertwining isomorphism.

An automorphism φ is called exponential if there exists a locally nilpotent derivation D such that

φ = exp(D) = D^{0} + D + \frac{1}{2!}D^{2} + · · · .

For instance, any automorphism of the form

φ(x_{n}) = x_{n} + f(x_{0}, x_{1}, …, x_{n−1}), f ∈ ℤ[x_{0}, x_{1}, …, x_{n−1}],

is exponential, see Drensky and Yu [10]. Nowicki [8, Proposition 6.1.4] shows that for any exponential automorphism φ = exp(D),

ℚ[x_{0}, x_{1}, …, x_{n}]^{φ} = ker D.  (1)

We introduce an analogue of these notions for the integer polynomial transformations.

Definition 3. The transformation D(F(A)) := (D(f_{n}(x_{0}, x_{1}, …, x_{m}))|_{(a_{0},a_{1},...,a_{m})})_{n≥0}, is called the D-derivative of the polynomial transformation F = (f_{n}(a_{0}, a_{1}, …, a_{m}))_{n≥0}, f ∈ ℤ[x_{0}, …, x_{m}].

Definition 4. A transformation F is called exponential if there exists a locally nilpotent derivation D such that

F(A) = exp D(A).

We can now rewrite Theorem 1 for an exponential transformation.

Theorem 5. Suppose a transformation F is exponential and F(A) = exp D(A) for some locally nilpotent derivation D. Then a polynomial transformation G is F-invariant if and only if D(G(A)) = 0, where 0 stands for the zero sequence (0, 0, 0, …).

Proof. Suppose that the transformation G is F-invariant. Then by Theorem 1 we have

G(A) = (g_{n}(a_{0}, a_{1}, …, a_{m}))_{n≥0} ,

where g_{n}(x_{0}, x_{1}, …, x_{m}) ∈ ℤ[x_{0}, x_{1}, …, x_{m}]^{φ}, for φ = exp D. Since the automorphism φ is exponential, we have that D(g_{n}(x_{0}, x_{1}, …, x_{m})) = 0 by (1). Thus D(G(A)) = 0.

Suppose now that the transformation G has the form

G(A) = (g_{n}(a_{0}, a_{1}, …, a_{m}))_{n≥0} ,

and D(G(A)) = 0. This implies that D(g_{n}(a_{0}, a_{1}, …, a_{m})) = 0 for all n. Then we have that the polynomial g_{n}(x_{0}, x_{1}, …, x_{m}) belongs to ℚ[x_{0}, x_{1}, …, x_{n}]^{φ} where φ = exp D. By Theorem 1 the transformation G is F-invariant.

The Weitzenbök derivations are related to some special transformations by the following theorem:

Theorem 6. Given an integer sequence (α_{n})_{n≥0}, the transformation F (A)= (a_{n} + ∑_{i=0}^{n−1} α_{i}a_{i})_{n≥0}

is exponential and F (A) = exp D(A), where the derivation D is a Weitzenbök derivation defined by

D(f) = ∑_{i=1}^{∞} \frac{(−1)^{i+1}}{i} E^{i}(f),

and E(x_{k}) = ∑_{i=0}^{k−1} α_{k}x_{k}.

The proof follows from van den Essen [9, Proposition 2.1.3].

Theorem 6 yields an algorithm to solve Problem 1 in the case when the transformation F(A) = (b_{n})_{n≥0} has the special form

b_{n} = a_{n} + ∑_{i=0}^{n−1} α_{i}a_{i}, α_{i} ∈ ℤ.

In this case, for the corresponding polynomial automorphism φ(x_{n}) = x_{n} + ∑_{i=0}^{n−1} α_{i}x_{i}, we find

the explicit form of the Weitzenbök derivation D such that φ = exp(D) (as predicted in Theorem 5). After that we find any (D, D)-intertwining automorphism Ψ and obtain that ker D = Ψ (ker D) . Then an arbitrary sequence of kernel elements defines an F-invariant transformation (by Theorem 1).

To solve Problem 2 for a transformation F(A) = (b_{n} = f_{n}(a_{0}, a_{1}, …, a_{m}))_{n≥0} we find a locally nilpotent derivation D of ℤ[x_{0}, x_{1}, …, x_{m}] such that f_{n}(x_{0}, x_{1}, …, x_{m}) ∈ ker D. This can be done by the method of undetermined coefficients. We define the automorphism φ = exp D and the transformation G(A) = (b_{n} = φ(a_{n}))_{n≥0}. Then the transformation F is G-invariant by Theorem 1.

3 The μ-binomial transformations

We use the techniques developed in Section 2 to get another proof of the following well known result.

Theorem 7 ([1, 2]). The Hankel transformation H is B_{μ}-invariant.

Proof. We follow the algorithm from Section 2. The automorphism φ_{μ} corresponding to B_{μ} has the form:

φ_{μ}(x_{n}) = ∑_{i=0}^{n} \binom{n}{i}x_{i}μ^{n−i}.

For the basic Weitzenbök derivation D we have

exp(μD)(x_{n}) = ∑_{i≥0} \frac{1}{i!} (μD)^{i} (x_{n}) = ∑_{i=0}^{n} \frac{n(n−1) ⋯ (n−(i−1))}{i!} μ^{i}x_{n−i} =

= ∑_{i=0}^{n} \binom{n}{i}μ^{i}x_{n−i} = ∑_{i=0}^{n} \binom{n}{i}x_{i}μ^{n−i}.

Thus φ_{μ} = exp(μD). It follows that the transformation B_{μ} is exponential, i.e., B_{μ} = exp(μD)A. Since the catalectiant belongs to the kernel of the derivation D we have that D(H(A)) = 0. Then by Theorem 5 we obtain that the transformation H is B_{μ}-invariant.

The map exp(μD) : ℚ[x_{0}, x_{1}, …, x_{n}] → ℚ[x_{0}, x_{1}, …, x_{n}] is a ring homomorphism, see van den Essen [9, Proposition 1.2.24]. It follows that φ_{μ_{1}+μ_{2}} = φ_{μ_{1}} ∘ φ_{μ_{2}}. Therefore φ_{μ} ∘ φ_{−μ} is the identity map and B_{μ}^{−1} = B_{−μ}. It follows immediately that the inverse transformation B_{μ}^{−1} is also an H-invariant transformation.

See French [11] for a proof that all B_{μ}-invariant transformations form a group. The identity φ_{μ_{1}+μ_{2}} = φ_{μ_{1}} ∘ φ_{μ_{2}} implies that the group (ℤ, +) is a subgroup of those groups.

The following theorem gives a solution to Problem 1 for the μ-binomial transformation.

Theorem 8. A transformation F is B_{μ}-invariant if and only if D (F(A)) = 0.

The proof follows from Theorem 5.

The next result follows from Bedratyuk [12, Theorem 3.2].

Theorem 9. Let F be an arbitrary B_{μ}-invariant transformation. Then F(1) = 0, where 1 = (1, 1, 1, …, 1, …).

Below we describe some Hankel-type transformations which arise in classical invariant theory. Note that all of these transformations are B_{μ}-invariant and B_{μ}^{−1}-invariant.

3.1 Cayley transformation

Put CAYLEY(A) = (b_{n+2})_{n≥0}, where

b_{n} = ∑_{i=0}^{n−2}(−1)^{i}\binom{n}{i}a_{n−i}a_{1}^{i}a_{0}^{n−k−1} + (n − 1)(−1)^{n+1}a_{1}^{n}.

The definition of this transformation is inspired by Theorem 2.

3.2 Transvectant transformation

Let A = (a_{n})_{n≥0} and C = (c_{n})_{n≥0} be two sequences. The transformation TR(A, C) = (b_{n})_{n≥0}, where

b_{n} = ∑_{i=0}^{n}(−1)^{i}\binom{n}{i}a_{i}c_{n−i},

is called the transvectant transformation. We have

Tr(B_{μ}(A), B_{μ}(C)) = Tr(A, C).

In the case C = A we get

b_{n} = ∑_{i=0}^{n}(−1)^{i}\binom{n}{i}a_{i}a_{n−i}.

3.3 Resultant transformation

Let A = (a_{n})_{n≥0} and C = (c_{n})_{n≥0} be two sequences. The transformation RES(A, C) = (b_{n})_{n≥0}, where b_{n} is the leading coefficient of the resultant of the polynomials

P_{n}(A) = ∑_{i=0}^{n} \binom{n}{i}a_{i}X^{n−i}, P_{n}(C) = ∑_{i=0}^{n} \binom{n}{i}c_{i}X^{n−i},

is called the resultant transformation.

3.4 Discriminant transformation

The transformation DISCR(A) = (b_{n})_{n≥0}, where b_{n} is the discriminant of the polynomial

P_{n+2}(A) = \frac{1}{(n + 2)^{n+2}} ∑_{i=0}^{n+2} a_{i}\binom{n + 2}{i}X^{n+2−i},

is called the discriminant transformation. Problem 3. What is the explicit form of the (D, D)-intertwining isomorphism Ψ(F) for

F ∈ {CAYLEY, H, RES, DISCRIM, TR}?

4 Examples

4.1 Transformation PSUM(A) = (b_{n} = a_{0} + a_{1} + ⋯ + a_{n})_{n≥0}

The corresponding locally nilpotent derivation (see Theorem 5) has the form

D(x_{n}) = ∑_{i=1}^{∞} \frac{(−1)^{i+1}}{i} E^{i}(x_{n}).

We have

E(x_{0}) = 0, E(x_{n}) = x_{0} + x_{1} + x_{2} + ⋯ + x_{n−1},

E^{2}(x_{n}) = ∑_{i=0}^{n−1} E(x_{i}) = ∑_{i=0}^{n−1} ∑_{j=0}^{i−1} x_{j} = ∑_{i=0}^{n−2}(n − 1 − i)x_{i}.

By induction we obtain E^{i}(x_{n}) = ∑_{k=0}^{n−i} \binom{n − i − 1}{i − 1}x_{k}. Then

D(x_{n}) = ∑_{i=1}^{n} \frac{(−1)^{i+1}}{i} ∑_{k=0}^{n−i} \binom{n − i − 1}{i − 1}x_{k} = ∑_{k=0}^{n−1} ( ∑_{i=0}^{n−1−k} \frac{(−1)^{i}}{i + 1} \binom{n−1−k}{i}) x_{k} = ∑_{k=0}^{n−1} \frac{x_{k}}{n − k} .

Let us find a (D, D)-intertwining transformation Ψ. We show that

Ψ(A) = \{Ψ(x_{n}) = ∑_{k=0}^{n}(−1)^{n+k}k! \begin{cases}n \\ k \end{cases}\} x_{k}, Ψ(x_{0}) = x_{0}\} ,

where \begin{cases}n \\ k \end{cases}\} is the Stirling number of the second kind, is such a transformation. In fact,

D (Ψ(x_{n})) = D (∑_{k=0}^{n}(−1)^{n+k}k! \begin{cases}n \\ k \end{cases}\} x_{k}) = ∑_{k=0}^{n}(−1)^{n+k}k! \begin{cases}n \\ k \end{cases}\} ∑_{i=0}^{k−1} \frac{x_{i}}{k − i}

= ∑_{i=0}^{n−1} ∑_{j=i+1}^{n}(−1)^{n+j} \begin{cases}n \\ j \end{cases}\} \frac{j!}{j − i} x_{i} = n ∑_{i=0}^{n−1}(−1)^{n−1+i} \begin{cases}n − 1 \\ i \end{cases}\} i! x_{i} = Ψ(D(x_{n})).

Therefore, we may now construct a PSUM-invariant transformation using our known B_{μ}- invariant transformations and this (D, D)-intertwining transformation Ψ. For instance, the transformation

Ψ(H(A)) = {a_{0}, −a_{1}^{2} − a_{1}a_{0} + 2 a_{2}a_{0}, −4 a_{1}a_{2}a_{0} + 24 a_{1}a_{2}a_{3} + 24 a_{0}a_{1}a_{3} + 48 a_{0}a_{2}a_{4} − 8 a_{2}^{3}− −8 a_{0}a_{2}^{2} − 12 a_{1}a_{2}^{2} − 36 a_{0}a_{3}^{2} − 4 a_{1}^{2}a_{2} − 24 a_{1}^{2}a_{4} + 24 a_{1}^{2}a_{3} − 24 a_{0}a_{1}a_{4}, …},

is PSUM-invariant.

4.2 The Transformation SUM(A) = (b_{n} = a_{n} + a_{n−1})_{n≥0}

We have φ(x_{n}) = x_{n} + x_{n−1}, E(x_{n}) = φ(x_{n}) − x_{n} = x_{n−1}, and

D(x_{n}) = ∑_{i≥1} \frac{(−1)^{i+1}}{i} E^{i}(x_{n}) = ∑_{i=1}^{n} \frac{(−1)^{i+1}}{i} x_{n−i}.

Let

Ψ(x_{0}) = x_{0}, Ψ(x_{n}) = c_{n,1}x_{1} + c_{n,2}x_{2} + ⋯ + c_{n,n}x_{n}.

The (D, D)-intertwining map satisfies the conditions D(Ψ(x_{n})) = Ψ (D(x_{n})), n = 0, 1, 2, … n.

After a routine calculation we get that c_{n,i} = i! \begin{cases}n \\ i \end{cases}\} and a (D, D)-intertwining map is given

by

Ψ(x_{n}) = ∑_{i=1}^{n} i! \begin{cases}n \\ i \end{cases}\} x_{i}.

Thus, the transformation

Ψ(H(A)) = |\begin{matrix}Ψ(a_{0}) & & Ψ(a_{1}) & & Ψ(a_{2}) & & ⋯ & & & Ψ(a_{n}) \\ Ψ(a_{1}) & & Ψ(a_{2}) & & Ψ(a_{3}) & & ⋯ & & & Ψ(a_{n+1}) \\ . & . & . & . & . & . & . & . & . & . & . & . & . & . & . & . & . & . & . & . & . & . & . & . & . & . & . & . & . & . & . & . & . & . & . & . & . & . & . & . & . & . \\ Ψ(a_{n−1}) & & Ψ(a_{n}) & & Ψ(a_{n−1}) & & ⋯ & & & Ψ(a_{2n−1}) \\ Ψ(a_{n}) & & Ψ(a_{n+1}) & & Ψ(a_{n+2}) & & ⋯ & & & Ψ(a_{2n}) \end{matrix}| .

is SUM-invariant.

4.3 Transformation DIFF(A) = (b_{n} = a_{n} − a_{n−1})_{n≥0}

The corresponding automorphism has the form φ(x_{n}) = x_{n}−x_{n−1}. This implies that E(x_{n}) = −x_{n−1} and E^{i}(x_{n}) = (−1)^{i}x_{n−i}. Then the derivation D is defined by

D(x_{n}) = ∑_{i≥1} \frac{(−1)^{i+1}}{i} E^{i}(x_{n}) = ∑_{i=1}^{n} \frac{(−1)^{i+1}}{i} (−1)^{i}x_{n−i} = − ∑_{i=1}^{n} \frac{x_{n−i}}{i} .

A (D, D)-intertwining map has the form

Ψ(x_{n}) = ∑_{i=1}^{n}(−1)^{i}i! \begin{cases}n \\ i \end{cases}\} x_{i}.

4.4 The transformation F = (b_{n} = ∑_{i=0}^{2n}(−1)^{i}a_{i}a_{2n−i})_{n≥0}.

Let us try to solve Problem 2 for this transformation. To do so, we need to find a suitable locally nilpotent derivation that satisfies the conditions

D (∑_{i=0}^{2n}(−1)^{i}x_{i}x_{2n−i}) = 0, n = 0, 1, ….

Let us consider the locally nilpotent derivation D with D(x_{i}) = x_{i−1}. Now

D (∑_{i=0}^{2n}(−1)^{i}x_{i}x_{2n−i}) = 0,

as is shown in Bedratyuk [13]. Let us calculate the exponential automorphism φ = exp D. We have

φ(x_{n}) = D^{0}(x_{n}) + D(x_{n}) + \frac{1}{2!}D^{2}(x_{n}) + ⋯ = x_{n} + x_{n−1} + \frac{1}{2!} x_{n−2} + \frac{1}{n!} x_{0}.

Define a rational transformation by G(A) := (φ(x_{n}))_{n≥0}. Then F(G(A)) = F(A).

5 Acknowledgments

The author would like to thank the referees for many valuable suggestions that improved the paper.

References

[1] J. Layman, The Hankel transform and some of its properties, J. Integer Seq., 4 (2001), Article 01.1.5.

[2] M. Spivey and L. Steil, The k-binomial transforms and the Hankel transform, J. Integer Seq., 9 (2006), Article 06.1.1.

[3] O. Glenn, Treatise on Theory of Invariants, Boston, 1915.

[4] J. Grace and J. A. Young, The Algebra of Invariants, Cambrige Univ. Press, 1903.

[5] J. J. Sylvester, On the principles of the calculus of forms, Cambridge and Dublin Math. J., 7 (1852), 52–97.

[6] S. Lie, Theorie der Transformationsgruppen. Erster Abschnitt, Teubner, 1888.

[7] D. Hilbert, Theory of Algebraic Invariants, Cambridge University Press, 1993.

[8] A. Nowicki, Polynomial Derivations and their Rings of Constants, UMK, Torun, 1994.

[9] A. van den Essen, Polynomial Automorphisms and the Jacobian Conjecture, Birkhäuser, 2000.

[10] V. Drensky and J.-T. Yu, Exponential automorphisms of polynomial algebras, Comm. Algebra 26 (1998), 2977–2985.

[11] C. French, Transformations preserving the Hankel transform, J. Integer Seq., 10 (2007), Article 07.7.3.

[12] L. Bedratyuk, Semi-invariants of binary forms and identities for Bernoulli, Euler and Hermite polynomials, Acta Arith., 151 (2012), 361–376.

[13] L. Bedratyuk, Kernels of derivations of polynomial rings and Casimir elements, Ukr. Mat. Zh., 62 (2010), 435–452.

2010 Mathematics Subject Classification: Primary 11B75; Secondary 13A50. Keywords: invariant, polynomial transformation, Hankel determinant.

Received March 3 2012; revised versions received June 26 2012; August 6 2012; August 23 2012. Published in Journal of Integer Sequences, September 8 2012. Return to Journal of Integer Sequences home page.